# An Embedding is Worth a Thousand Noisy Labels

### v1

@@ Abstract
The performance of deep neural networks scales with dataset size and label quality, rendering the efficient mitigation of low-quality data annotations crucial for building robust and cost-effective systems. Existing strategies to address label noise exhibit severe limitations due to computational complexity and application dependency. In this work, we propose WANN, a Weighted Adaptive Nearest Neighbor approach that builds on self-supervised feature representations obtained from foundation models. To guide the weighted voting scheme, we introduce a reliability score $\eta$, which measures the likelihood of a data label being correct. WANN outperforms reference methods, including a linear layer trained with robust loss functions, on diverse datasets of varying size and under various noise types and severities. WANN also exhibits superior generalization on imbalanced data compared to both Adaptive-NNs (ANN) and fixed $k$-NNs. Furthermore, the proposed weighting scheme enhances supervised dimensionality reduction under noisy labels. This yields a significant boost in classification performance with 10×10$\times$10 × and 100×100$\times$100 × smaller image embeddings, minimizing latency and storage requirements. Our approach, emphasizing efficiency and explainability, emerges as a simple, robust solution to overcome the inherent limitations of deep neural network training. The code is available at github.com/francescodisalvo05/wann-noisy-labels.
keywords:
Label noise, Efficiency, k-NN, Adaptive-NN, Weighted-NN, Robust loss functions
Figure 1: Illustration of the proposed Weighted Adaptive Nearest Neighbor (WANN) algorithm. Initially, a reliability score ($\eta$) is computed for each training observation, representing the inverse of the minimum number of samples needed for a correct prediction. During inference, the adaptive neighborhood size ($k_{T}$) of each test observation ($x_{T}$) is determined based on the reliability score of its closest training point ($k_{T}=\frac{1}{\eta}=3$). A weighted majority vote determines the final label, reducing the impact of noisy labels.
@@ 1 Introduction
The remarkable results achieved by deep neural networks in a multitude of domains are possible due to ever-increasing computational power. This progress has enabled the development of deeper architectures with stronger learning capabilities. However, these high-parametric architectures are often characterized as data-hungry, because they require a large amount of data to generalize effectively. Collecting and annotating such extensive datasets can be both expensive and time-consuming, potentially introducing machine and human errors. In fact, the proportion of corrupted labels in real-world datasets ranges from 8-38.5% [42]. While unsupervised and semi-supervised methods have garnered significant attention within the research community lately [7, 44, 1], supervised methods are still widely employed due to their generally higher performance. In this context, incorrect labels pose significant challenges, including a tendency to memorize label noise, negatively affecting generalization capabilities. This issue, as demonstrated by [56], is hindering the adoption of AI systems in safety-critical domains, such as healthcare.
Consequently, there is a growing interest in enhancing the robustness of deep models against noisy labels, resulting in five different lines of research [42]: robust architectures, robust regularization, robust loss function, loss adjustment and sample selection. Nevertheless, as also highlighted by [58], all these approaches involve a learning process and, by definition, have limitations in generalizing to diverse datasets or varying noise rates.
@@
We address those limitations and the challenges posed due to noisy labels by exploiting feature representations obtained from large pre-trained models, commonly referred to as foundation models. Many of these foundation models are now publicly available and designed for a wide range of applications, including image classification, semantic segmentation tasks, and beyond. While their initial focus was primarily on text [12, 6] and natural images [11, 9, 34, 38], there is currently a focus on the development of open-source foundation models tailored to specific domains, such as healthcare [43, 29, 57, 52], making it possible to translate this paradigm to various applications with little effort. Specifically, large image-based foundation models are usually trained in a self-supervised fashion, e.g., by means of contrastive learning [7], representing similar objects closely within the embedding space, and semantically distinct objects further away. Thus, the position of an image in the embedding space can be just as informative as the label, if not more so. Although embedding-space approaches are not new in the literature to process noisy labels, they typically focus on noise detection, through online [5] or offline [58] methods. However, thanks to the representational power of modern embedding spaces [38, 34], we believe that it is also possible to provide robust predictions by building on simple $k$-NN methods. This category of approaches not only offers computational efficiency but also enhances explainability. Moreover, the lower reliance on hyperparameters further enhances generalizability. Thus, owing to its simplicity and efficiency, it is able to mitigate some of the limitations associated with the training of deep networks.
In this work, we propose a Weighted Adaptive Nearest Neighbor (WANN) approach, illustrated in Figure 1, which operates on the image embeddings extracted from a pre-trained foundation model. The introduced weighted adaptive voting scheme addresses the challenges posed by large-scale, limited, and imbalanced noisy datasets. Our contributions are:
@@
Consequently, through lightweight embeddings and a simple yet efficient classification algorithm, we demonstrate that working efficiently within the embedding space might constitute a potential paradigm shift, increasing efficiency and robustness while addressing several critical limitations associated with training deep neural networks.
@@
Figure 2: Example of test images and their relative top-$3$ closest training samples extracted from the STL-10 dataset [8]. Figure 2(a) shows two test images having wrong labels, while Figure 2(b) shows ambiguous labels, including multiple known objects within a single image. For instance, a bird with a ship in the background, or a dog fighting with a cat.
2 Related works
Noise robust learning   Robust learning from noisy labels has received considerable attention in recent years, and multiple deep-learning-based methods are currently available [42]. Effective and popular approaches involve multi-network learning, where noisy samples during training are discarded. The sample selection might be guided by a mentor network [23] or a peer network [19, 53, 46]. However, due to their computational complexity, these models often lack generalizability across datasets or noise settings. Some of the lightweight techniques designed to handle noisy labels fall under the category of robust loss functions. While the Cross Entropy (CE) suffers from overfitting to wrong labels [56], the Mean Absolute Error (MAE) has been theoretically guaranteed to be noise-tolerant [15]. However, it suffers from severe underfitting in challenging domains. To address this issue and enhance its generalization, the Generalized Cross Entropy (GCE) [60] was introduced as a generalization of both MAE and CE. While [47] propose the Symmetric Cross Entropy (SCE), [59] overcome the symmetric condition through Asymmetric Loss Functions. Among state-of-the-art methods, we recall Early-Learning Regularization [28] which tries to prevent memorization of the false labels via regularization, and Active Passive Losses (APL) [31] which combines two robust losses to balance between overfitting and underfitting. Recently, Active Negative Losses (ANL) [54] replaced the previous passive loss (in APL) with a Normalized Negative Loss Function (NNLF). Nevertheless, despite their relatively lower computational complexity compared to multi-network methods, they still inherit challenges related to training neural networks. Some of these challenges include the demand for large datasets, lack of explainability, computational complexity, hyperparameter dependence, and overfitting towards wrong labels.
$k$-NN for label noise   Besides native deep learning-based approaches, certain traditional machine learning methods are robust by design. Notably, $k$-Nearest-Neighbors-based methods have recently gained attention for their utility in filtering out noisy training observations [37, 5, 26]. Although the proposed online $k$-NN approaches can enhance the robustness of DNNs trained with label noise, it does not address the fundamental issues associated with DNNs, such as the lack of explainability, data-efficiency, and generalizability. A closely related work [58] already emphasized the potential of a training-free cleaning strategy based on $k$-NN, relying on CLIP [38] embeddings. In alignment with this research direction, we aim to accentuate the intrinsic robustness and explainability benefits (cf. Figure 2) of $k$-NN in classification tasks. This is further enhanced by the adoption of an adaptive neighborhood for each test observation, defined through a reliability scoring.
@@ 3 Method
At its core, the proposed method relies on an adaptive $k$-NN search over the image embeddings extracted from a pre-trained foundation model, using the Euclidean distance as reference distance metric. In particular, we employ the DINOv2 [34] Large backbone, with $14{\times}14$ patches, yielding image embeddings of size $1024$. DINOv2 was trained in a self-supervised fashion on $142$M images, and the choice of the backbone is further motivated in Section 4.1. To enhance computational efficiency, we pre-generated a database of embeddings for all of the evaluated datasets, following the approach adopted by [33].
@@
Label reliability score   Although $k$-NN is robust against noisy labels by design, its efficacy may diminish under severe noise conditions. To address this challenge, the introduction of a weighting scheme becomes helpful, aiming to mitigate the impact of noisy labels. Intuitively, a proper weighting scheme should assign higher weights to samples that are more likely to have a correct label, and lower weights to those potentially mislabeled. In this context, we introduce the concept of label reliability, building upon the work of [40]. Extending their approach, which determined an optimal $k$ value for each training observation, we determine the reliability (i.e., quality) of the associated label, guiding a subsequent weighting scheme. For a given training observation $x_{i}$ and its associated label $y_{i}$, the reliability $\eta_{i}$ is defined as the inverse of the minimum number of training samples (neighborhood size) required for a correct $k$-NN classification. Dropping the index $i$ for readability, we express $\eta$ as the function $\mathcal{H}(x,y;\mathcal{X})$, where $\mathcal{X}$ represents the training dataset:
$\mathcal{H}(x,y;\mathcal{X})=\frac{1}{\underset{{k\in[k_{\text{min}},k_{\text{max}}]}}{\arg\,\min}{\,}{\texttt{kNN}_{\mathcal{X}}(x,k)=y}}$ (1)
@@
Here $k_{\text{min}}$ and $k_{\text{max}}$ define the search interval for $k\in\mathbb{N}$. It was previously shown by [58] that for a fixed $k$-NN approach a suitable value for $k$ is $10$ to detect corrupted labels. Consequently, we set $k_{\text{min}}=11$ and $k_{\text{max}}=51$. This allows to establish a substantial search space covering a wide range of noise levels while ensuring a baseline ability of detecting corrupt labels ($k>10$). During parameter search we increase $k$ by $2$ at each incremental step, ensuring an odd number of samples in the neighborhood while reducing the computational complexity.
@@
The correlation between the value of $k$ and the reliability of the assigned label stems from the main property of the embedding space. The proximity of similar objects suggests that a small $k$ should be sufficient for accurate classification. In such cases, the reliability of the data point is higher ($k\downarrow\;\eta\uparrow$). Conversely, if a data point requires a large $k$ value for correct classification, approaching or reaching $k_{\text{max}}$, it may signal potential mislabeling ($k\uparrow\;\eta\downarrow$). The pseudocode is presented in Algorithm 1, and it requires three parameters: $k_{\text{min}},k_{\text{max}}$, and $D_{\text{train}}$. The latter, $D_{\text{train}}$, represents a data structure containing a triple $(\texttt{key},x,y)$ for each train observation. Here, key serves as a unique identifier.
@@
  $P\leftarrow\texttt{PairwiseDist}(D_{\text{train}},D_{\text{train}})$matrix}
@@
     for $k=k_{\text{min}},k_{\text{min}}+2,\ldots,k_{\text{max}}$ do
        $\hat{y}\leftarrow\texttt{kNN}_{P}(x,k)$$k$ points excluding $x$}
@@
           $R[\texttt{key}]\leftarrow\frac{1}{k}$
@@
Weighted Adaptive Nearest Neighbor (WANN)   Adaptive Nearest Neighbors algorithms were commonly employed for tabular data with limited number of observations and classes [16, 45, 48]. Following the intuition of [40], we use $\eta$ to adaptively determine the neighborhood size for any test observation. In fact, if a test sample is close to a potentially noisy label ($\eta\downarrow$) a larger neighborhood is preferable to reduce variance. Conversely, if a test observation is close to a possibly clean label ($\eta\uparrow$), a small neighborhood increases specificity. Thus, given a test observation $x_{T}$ with nearest training sample $(x_{n},y_{n},\eta_{n})$ (nearest in the embedding space), its test neighborhood size $k_{T}$ will be adaptively defined by means of the reliability score of $x_{n}$ (cf. Equation 2).
@@
With this notation we formulate an Adaptive Nearest Neighbor (ANN) method as baseline, similar to [40]:
@@
To address this limitation, we employ WANN’s reliability score $\eta$ to filter out potentially wrong labels (i.e., $\eta=1/{k_{max}}$) before performing LDA. We term this approach Filtered LDA (FLDA). Once the projection axes have been obtained through the clean dataset, we subsequently project both training and test datasets onto those axes.

### v3

@@ Abstract
The performance of deep neural networks scales with dataset size and label quality, rendering the efficient mitigation of low-quality data annotations crucial for building robust and cost-effective systems. Existing strategies to address label noise exhibit severe limitations due to computational complexity and application dependency. In this work, we propose WANN, a Weighted Adaptive Nearest Neighbor approach that builds on self-supervised feature representations obtained from foundation models. To guide the weighted voting scheme, we introduce a reliability score $\eta$, which measures the likelihood of a data label being correct. WANN outperforms reference methods, including a linear layer trained with robust loss functions, on diverse datasets of varying size and under various noise types and severities. WANN also exhibits superior generalization on imbalanced data compared to both Adaptive-NNs (ANN) and fixed $k$-NNs. Furthermore, the proposed weighting scheme enhances supervised dimensionality reduction under noisy labels. This yields a significant boost in classification performance with 10×10$\times$10 × and 100×100$\times$100 × smaller image embeddings, minimizing latency and storage requirements. Our approach, emphasizing efficiency and explainability, emerges as a simple, robust solution to overcome inherent limitations of deep neural network training. The code is available at github.com/francescodisalvo05/wann-noisy-labels.
@@ 1 Introduction
The remarkable results achieved by deep neural networks in a multitude of domains are possible due to ever-increasing computational power. This progress has enabled the development of deeper architectures with stronger learning capabilities. However, these high-parametric architectures are often characterized as data-hungry, because they require a large amount of data to generalize effectively. Collecting and annotating such extensive datasets can be both expensive and time-consuming, potentially introducing machine and human errors. In fact, the proportion of corrupted labels in real-world datasets ranges from $8\%$ to $38.5\%$ (Song et al., 2022). While unsupervised and semi-supervised methods have garnered significant attention within the research community (Chen et al., 2020; Tarvainen & Valpola, 2017; Berthelot et al., 2019), supervised methods are still widely employed due to their generally higher performance. In this context, incorrect labels pose significant challenges, including a tendency to memorize label noise, negatively affecting generalization capabilities. This issue, as demonstrated by Zhang et al. (2021), is hindering the adoption of AI systems in safety-critical domains, such as healthcare. Consequently, there is a growing interest in enhancing the robustness of deep models against noisy labels, resulting in five different lines of research (Song et al., 2022): robust architectures, robust regularization, robust loss function, loss adjustment and sample selection. Nevertheless, as also highlighted by Zhu et al. (2022), all these approaches involve a learning process and, by definition, have limitations in generalizing to diverse datasets or varying noise rates. We address those limitations and the challenges posed due to noisy labels by exploiting feature representations obtained from large pre-trained models, commonly referred to as foundation models. Many of these foundation models are now publicly available and designed for a wide range of applications, including image classification, semantic segmentation tasks, and beyond. While their initial focus was primarily on text (Devlin et al., 2019; Brown et al., 2020) and natural images (Dosovitskiy et al., 2021; Caron et al., 2021; Oquab et al., 2023; Radford et al., 2021), there is currently a focus on the development of open-source foundation models tailored to specific domains, such as healthcare (Tu et al., 2024; Li et al., 2023; Zhou et al., 2023; Xu et al., 2024), making it possible to translate this paradigm to various applications with little effort. Specifically, large image-based foundation models are usually trained in a self-supervised fashion, e.g., by means of contrastive learning (Chen et al., 2020), representing similar objects closely within the embedding space, and semantically distinct objects further away. Thus, the position of an image in the embedding space can be just as informative as the label, if not more so. Although embedding-space approaches are not new in the literature to process noisy labels, they typically focus on noise detection, through online (Bahri et al., 2020) or offline (Zhu et al., 2022) methods. However, thanks to the representational power of modern embedding spaces (Radford et al., 2021; Oquab et al., 2023), we believe that it is also possible to provide robust predictions by building on simple $k$-NN methods. This category of approaches not only offers computational efficiency but also enhances explainability. Moreover, the lower reliance on hyperparameters further enhances generalizability. Thus, owing to its simplicity and efficiency, it is able to mitigate some of the limitations associated with the training of deep networks. In this work, we propose a Weighted Adaptive Nearest Neighbor (WANN) approach, illustrated in Figure 1, which operates on the image embeddings extracted from a pre-trained foundation model. The introduced weighted adaptive voting scheme addresses the challenges posed by large-scale, limited, and imbalanced noisy datasets. Our contributions are:
@@
Consequently, through lightweight embeddings and a simple yet efficient classification algorithm, we demonstrate that working efficiently within the embedding space might constitute a potential paradigm shift, increasing efficiency and robustness while addressing critical limitations related to model training.
Figure 1: Illustration of the proposed Weighted Adaptive Nearest Neighbor (WANN) algorithm. Initially, a reliability score ($\eta$) is computed for each training observation, representing the inverse of the minimum number of samples needed for a correct prediction. During inference, the adaptive neighborhood size ($k_{T}$) of each test observation ($x_{T}$) is determined based on the reliability score of its closest training sample ($k_{T}=\frac{1}{\eta}=3$). A weighted majority vote determines the final label, reducing the impact of noisy labels.
2 Related works
Noise robust learning   Robust learning from noisy labels has received considerable attention in recent years, and multiple deep-learning-based methods are currently available (Song et al., 2022). Two popular categories of approaches are sample selection and loss adjustment. Methods based on sample selection aim to isolate correctly labelled samples from the noisy dataset, often employing strategies such as multi-network learning, iterative data filtering, or both (Han et al., 2018; Jiang et al., 2018; Song et al., 2019; Li et al., 2020; Wei et al., 2020; Xiao et al., 2023; Zhang et al., 2024; Fooladgar et al., 2024). In contrast, loss adjustment techniques modify how the loss is computed for every sample, thereby mitigating the impact of label noise (Patrini et al., 2017; Tanaka et al., 2018; Liu et al., 2020; Zheng et al., 2021). However, due to their computational complexity, these models often lack generalizability across datasets or noise settings. Some of the lightweight techniques designed to handle noisy labels fall under the category of robust loss functions. While the Cross Entropy (CE) suffers from overfitting to wrong labels (Zhang et al., 2021), the Mean Absolute Error (MAE) has been theoretically guaranteed to be noise-tolerant (Ghosh et al., 2017). However, it suffers from severe underfitting in challenging domains. To address this issue and enhance its generalization, the Generalized Cross Entropy (GCE) (Zhang & Sabuncu, 2018) was introduced as a generalization of both MAE and CE. While Wang et al. (2019) propose the Symmetric Cross Entropy (SCE), Zhou et al. (2021) overcome the symmetric condition through Asymmetric Loss Functions. Active Passive Losses (APL) (Ma et al., 2020) which combines two robust losses to balance between overfitting and underfitting. Recently, Active Negative Losses (ANL) (Ye et al., 2023; 2024) replaced the previous passive loss (in APL) with a Normalized Negative Loss Function (NNLF). Nevertheless, despite their relatively lower computational complexity compared to multi-network methods, they still inherit challenges related to training neural networks. Some of these challenges include the demand for large datasets, lack of explainability, computational complexity, hyperparameter dependence, and overfitting towards wrong labels.
$k$-NN for label noise   Besides native deep learning-based approaches, certain traditional machine learning methods are robust by design. Notably, $k$-Nearest-Neighbors-based methods have recently gained attention for their utility in filtering out noisy training observations (Reeve & Kabán, 2019; Bahri et al., 2020; Kong et al., 2020). Although the proposed online $k$-NN approaches can enhance the robustness of DNNs trained with label noise, they do not address the fundamental issues associated with DNNs, such as the lack of explainability, data-efficiency, and generalizability. A closely related work (Zhu et al., 2022) already emphasized the potential of a training-free cleaning strategy based on $k$-NN with CLIP (Radford et al., 2021) embeddings, enhancing DNNs downstream performance. Furthermore, a concurrent work (Zhu et al., 2024) demonstrated that a simple noise-agnostic linear probing on high-quality features outperforms a deep network trained from scratch, achieving state-of-the-art results. In contrast, we show that a training-free approach can outperform linear probing, with or without robust loss functions. Indeed, our work represents a broader paradigm shift: rather than training large networks from scratch, we utilize high-quality feature representations from foundation models and employ a simple, interpretable, and training-free classification scheme in the embedding space. This design choice directly addresses practical challenges such as mitigating overfitting to noisy labels, enabling transparent predictions (cf. Figure 2), and simplifying dataset updates. We further enhance classification performance under label noise and label imbalance through the adoption of an adaptive neighborhood for each test observation.
@@ 2 Related works
Figure 2: Example of test images and their relative top-$3$ closest training samples extracted from the STL-10 dataset (Coates et al., 2011). Figure 2(a) shows two test images having wrong labels, while Figure 2(b) shows ambiguous labels, including multiple known objects within a single image. For instance, a bird with a ship in the background, or a dog fighting with a cat.
@@ 3 Method
At its core, the proposed method relies on an adaptive $k$-NN search over the image embeddings extracted from a pre-trained foundation model, using the Euclidean distance as reference distance metric. In particular, we employ the DINOv2 (Oquab et al., 2023) Large backbone, with $14{\times}14$ patches, yielding image embeddings of size $1024$. DINOv2 was trained in a self-supervised fashion on $142$M images, and the choice of the backbone is further motivated in Section 4.1. To enhance computational efficiency, we pre-generated a database of embeddings for all of the evaluated datasets, following the approach adopted by Nakata et al. (2022).
@@
Label reliability score   Although $k$-NN is robust against noisy labels by design, its efficacy may diminish under severe noise conditions. To address this challenge, the introduction of a weighting scheme becomes helpful, aiming to mitigate the impact of noisy labels. Intuitively, a proper weighting scheme should assign higher weights to samples that are more likely to have a correct label, and lower weights to those potentially mislabeled. In this context, we introduce the concept of label reliability, building upon the work of Sun & Huang (2010). Extending their approach, which inferred an optimal $k$ value for each training observation, we determine the reliability (i.e., quality) of the associated label, guiding a subsequent weighting scheme. For a given training observation $(x_{i},y_{i})$, we define the reliability $\eta_{i}$ as the inverse of the minimum number of training samples (neighborhood size) required for a correct $k$-NN classification. Dropping the index $i$ for readability, we express $\eta$ as the function $\mathcal{H}(x,y;\mathcal{X})$, where $\mathcal{X}$ represents the training dataset:
$\mathcal{H}(x,y;\mathcal{X})=\frac{1}{\underset{{k^{\prime}\in[k_{\text{min}},k_{\text{max}}]}}{\arg\,\min}{\,}{\texttt{kNN}_{\mathcal{X}}(x,k^{\prime})=y}}$ (1)
@@
Here $k_{\text{min}}$ and $k_{\text{max}}$ define the search interval for $k^{\prime}\in\mathbb{N}$. It was previously shown by Zhu et al. (2022) that for a fixed $k$-NN approach a suitable neighborhood size to detect corrupted labels is $10$. Consequently, we set $k_{\text{min}}=11$ and $k_{\text{max}}=51$. This allows to establish a substantial search space covering a wide range of noise levels while ensuring a baseline ability to detect corrupted labels ($k_{\text{min}}>10$). During parameter search we increase $k^{\prime}$ by $2$ at each incremental step, ensuring an odd number of samples in the neighborhood while reducing the computational complexity. If there is no $k^{\prime}$ satisfying Equation 1, we set the optimal $k=\frac{1}{k_{\text{max}}}$. The correlation between the value of $k$ and the reliability of the assigned label stems from the main property of the embedding space. The proximity of similar objects suggests that a small $k$ should be sufficient for accurate classification. In such cases, the reliability of the data point is higher ($k\downarrow\;\eta\uparrow$).
@@
Conversely, if a data point requires a large $k$ value for correct classification, approaching or reaching $k_{\text{max}}$, it may signal potential mislabeling ($k\uparrow\;\eta\downarrow$). The pseudocode is presented in Algorithm 1, and it requires three parameters: $k_{\text{min}},k_{\text{max}}$, and $D_{\text{train}}$. The latter, $D_{\text{train}}$, represents a data structure containing a triple $(\texttt{key},x,y)$ for each train observation. Here, key serves as a unique identifier.
@@
$P\leftarrow\texttt{PairwiseDist}(D_{\text{train}},D_{\text{train}})$triangleright matrix
@@
     for $k^{\prime}=k_{\text{min}},k_{\text{min}}+2,\ldots,k_{\text{max}}$ do
         $\hat{y}\leftarrow\texttt{kNN}_{P}(x,k^{\prime})$triangleright $k$ points excluding $x$
@@
              $R[\texttt{key}]\leftarrow\frac{1}{k^{\prime}}$
@@
Weighted Adaptive Nearest Neighbor (WANN)   Adaptive Nearest Neighbors algorithms were commonly employed for tabular data with a limited number of observations and classes (Geva & Sitte, 1991; Wettschereck & Dietterich, 1993; Wang et al., 2006). Following the intuition of Sun & Huang (2010), we use $\eta$ to adaptively determine the neighborhood size for any test observation. In fact, if a test sample is close to a potentially noisy label ($\eta\downarrow$) a larger neighborhood is preferable to reduce variance. Conversely, if a test observation is close to a possibly clean label ($\eta\uparrow$), a small neighborhood increases specificity. Thus, given a test observation $x_{T}$ with nearest training sample $(x_{n},y_{n},\eta_{n})$ (nearest in the embedding space), its test neighborhood size $k_{T}$ will be adaptively defined by means of the reliability score of $x_{n}$ (cf. Equation 2).
@@
With this notation we formulate an Adaptive Nearest Neighbor (ANN) method as baseline, similar to Sun & Huang (2010):
@@
While PCA seeks directions of maximum variance, LDA leverages class labels to maximize the separation between classes and minimize variance within classes, according to the Fisher-Rao criterion. Due to its supervised nature, LDA relies on correct labels, as label noise can result in inaccurate linear mappings. To address this limitation, we employ WANN’s reliability score $\eta$ to filter out potentially wrong labels (i.e., $\eta=1/{k_{max}}$) before performing LDA. We term this approach Filtered LDA (FLDA). Once the projection axes have been obtained through the clean dataset, we subsequently project both training and test datasets onto those axes.